OpenStudy (anonymous):

Are the ratios of 3 to 4 and 4 to 3 the same?

OpenStudy (mertsj):

Does 3/4 = 4/3?

OpenStudy (anonymous):

No

OpenStudy (ybarrap):

is 3/4 = 4/3 ? or is 3/4 - 4/3 = 0 or is (3*3 - 4*4)/(3*4) = 0 ?

OpenStudy (mertsj):

Then those two ratios are not the same.

OpenStudy (mathstudent55):

\( \dfrac{3}{4} = 0.75 \) \( \dfrac{4}{3} = 1.333... \) Answer: no.

OpenStudy (anonymous):

@Mertsj can you explain how to do ratios for me?

OpenStudy (mertsj):

I don't know what you want to know...a ratio is a fraction...it compares two quantities.

OpenStudy (ybarrap):

If you get 3 questions right out of a test that has 10 questions, you get 3/10 or .30. So your score would be 30% right

OpenStudy (mertsj):

It can be expressed in common fraction form or decimal fraction form.

OpenStudy (mertsj):

Let's say it snows 8 days in February and doesn't snow 20 days. The ratio of snow days to non-snow days is 8/20 The ratio of non-snow days to snow days is 20/8 The ratio of snow days to total days is 8/28 The ratio of non-snow days to total days is 20/28

OpenStudy (mertsj):

So you see in a ratio, the label is very important.

OpenStudy (mertsj):

Do you have a more specific question?

OpenStudy (mertsj):

A ratio can be written in 3 ways: 1. 3 to 4 2. 3/4 3. 3:4

OpenStudy (anonymous):

No I dont but that did help thank you :D

OpenStudy (mertsj):

yw